Show that


Ec[logƒ(Y)] = —— log(26²) —


7²+(m_µ)²


202


(c) Show that Kullback-Leibler divergence of f (y) with respect to g(y) is given by


D(glf) EG[logg (Y)]-EG[log f(Y)]


0² t² + (m-μ)² - 1}


1
Expert's answer
2021-05-07T14:39:47-0400

The Normal distribution of generating data g(y) and specified model f(y) is-

g(y)~N("m,t^2")

f(y)~N"(\mu,o^2)"



The probability density function is-


"Y=\dfrac{1}{\sqrt{2\pi}}e^{\frac{-x^2}{2}}"


"logY=log\dfrac{1}{\sqrt{2\pi}}+loge^{\frac{-x^2}{2}}=log\dfrac{1}{\sqrt{2\pi}}+\dfrac{-x^2}{2}~~~~~-(1)"



So "E(logY)=\int_{0}^{r}xlogYdx"


     "=\int_{0}^{r} xlog\dfrac{1}{\sqrt{2\pi}}+\dfrac{-x^3}{2}dx\\[9pt] =\dfrac{x^2}{2}log\dfrac{1}{\sqrt{2\pi}}|_0^r-\dfrac{x^4}{8}|_0^r\\[9pt]=\dfrac{r^2}{2}log\dfrac{1}{\sqrt{2\pi}}-\dfrac{r^4}{8}\\[9pt]=-log(2rt^2)-1"


Hence "E(logY)= -log(2rt^2)-1"